OpenStudy (okdutchman7):

Please help. I WILL MEDAL! @ganeshie8 @Hero @ParthKohli @Mehek14 @Kainui @TheSmartOne @Preetha @Directrix @mathmale @jhonyy9 @Zale101 @rebeccaxhawaii @Vocaloid @timo86m @calculusxy @Elsa213 @YanaSidlinskiy @Jaynator495

OpenStudy (okdutchman7):

A bus goes d miles in h hours and gets there 4 hours late. At what speed should the bus have been driven to have arrived on time?

OpenStudy (okdutchman7):

It's an abstract rate problem.

OpenStudy (anonymous):

4 miles per hour more

OpenStudy (anonymous):

to actually find out i need more details

OpenStudy (okdutchman7):

The answer is: \[\frac{ d }{ h-4 } \frac{ miles }{ hour }\]

OpenStudy (anonymous):

yeah

OpenStudy (okdutchman7):

I just don't know how to get there.

OpenStudy (anonymous):

v=d/t 4+v=(d/t)+4

OpenStudy (okdutchman7):

v is how fast it is going?

OpenStudy (anonymous):

yes

OpenStudy (okdutchman7):

I think I figured it out. Thank You!

OpenStudy (okdutchman7):

I have more, so I will post another question.

OpenStudy (anonymous):

make up your own numbers

OpenStudy (anonymous):

10 m/h in 6 hours yo go 60 m if you going at 4 mph in 4 hours you do 40 thus you will be 2 hours late. come up with formula